OpenStudy (anonymous):

transform the following polar equation into an equation in rectangular coordinates r=-4 sin theta would it be x^2(y+2)^2=4?

ganeshie8 (ganeshie8):

\(r = -4\sin\theta\) is a circle

ganeshie8 (ganeshie8):

the graph wont change between polar and cartesian

ganeshie8 (ganeshie8):

does your equation look anything like the eqn of a circle ?

OpenStudy (anonymous):

no so y=-4?

ganeshie8 (ganeshie8):

aren't you just guessing the options ?

OpenStudy (anonymous):

Remember \[x = r \cos \theta\]\[y= r \sin \theta\]\[r^2 = x^2+y^2 \implies r = \sqrt{x^2+y^2}\]

Parth (parthkohli):

Multiply both sides by \(r\) to see it.

OpenStudy (anonymous):

how would I plug in the numbers though?

OpenStudy (anonymous):

\[r^2 = -4rsin \theta \] doing as parth suggested, see what you can do with this given the information above.

ganeshie8 (ganeshie8):

\[\color{red}{r^2} = -4\color{blue}{r\sin\theta}\] you don't want to see \(r, \theta\) so replace \(\color{red}{r^2} \) by \(\color{red}{x^2+y^2} \) and \(\color{blue}{r\sin\theta}\) by \(\color{blue}{y}\) just algebra circus

OpenStudy (anonymous):

therefore X+Y=-4? I get it

OpenStudy (anonymous):

Not quite..

OpenStudy (anonymous):

\[x^2+y^2=-4y\] \[r^2 = x^2+y^2~~~~y = r \sin \theta\]

OpenStudy (anonymous):

And then from there you just substitute right?

ganeshie8 (ganeshie8):

technically we're done with the transformation \(x^2+y^2=-4y\) is the corresponding rectangular equation are you given options or something ?

OpenStudy (anonymous):

I was right actually because x^2(y+2)^2=4 was correct

ganeshie8 (ganeshie8):

you're correct upto a + sign

ganeshie8 (ganeshie8):

completing the square gives you \(x^2\color{red}{+}(y+2)^2=4\) which is slightly different from \(x^2(y+2)^2=4\) lexically, but totally different semantically. you might think "its just a +"... graph each of them and see

OpenStudy (jhannybean):

\[x^2+y^2=-4y\]\[x^2+(y^2+4y)=0\]\[x^2+(y^2+4y+\color{red}{4})=4\]\[\boxed{x^2+(y+2)^2 = 4}\]

OpenStudy (triciaal):